OpenStudy (anonymous):

Factor the polynomial completely: 4x^7-64x^3 So far i have (4x^3)(x^4-16) Looks really simple to finish, could someone help?

OpenStudy (anonymous):

x^4 - 16 can be simplified to (x-2) (x^3 + 2x^2 + 4x + 8)

OpenStudy (anonymous):

(4x^3)(x-2) (x^3 + 2x^2 + 4x + 8)

OpenStudy (anonymous):

It can be simplified more than that. You have that \[x^4-16=(x^2-4)(x^2+4)=(x-2)(x+2)(x^2+4)\] Thus, the fully simplified form is \[4x^3(x-2)(x+2)(x^2+4)\]